OpenStudy (anonymous):

Find the equation of the tangent line for the given function at x =(1/6) pi f(x)= 4sec(x)

OpenStudy (anonymous):

f'(x)= 4sec(x)tan(x) then...

OpenStudy (zzr0ck3r):

now plug in pi/6 this will give you the slope of the line and a point (pi/6,f(pi/6)), use this information and the form y = mx+b

OpenStudy (anonymous):

I dont understand what happens at sec(pi/6) or tan(pi/6)

OpenStudy (zzr0ck3r):

sec(pi/6) = 1/cos(pi/6) = 1/(sqrt(3)/2) = 2/sqrt(3)

OpenStudy (anonymous):

so the slope is 2?